Mr. Rahul has been saving Rs.9,500 at the beginning of every quarter for 25 years at an
average return of 14% p.a. compounded quarterly. He has to set apart Rs.15,00,000 for
his daughter’s marriage and Rs.10,00,000 for his son’s marriage from accumulated sum
after 25 years. He is saving separately for their education and he wishes to withdraw a
certain amount for the next 25 years on the remaining sum. The money accumulated will
be invested at a rate of 8% p.a. How much can he withdraw at the beginning of every
year?
1
Expert's answer
2020-12-10T13:42:42-0500

Sum of saving after 25 years:

"9.500*\frac{(1+\frac{0.14}{4})*((1+\frac{0.14}{4})^{25*4}-1)}{(1+\frac{0.14}{4})-1}=8,481,629.11"

He needs:

"15,000,000+10,000,000-8,481,629.11=16,518,370.89"

He needs to withdraw ever year:

"16,518,370.89=P*\frac{1.08*(1.08^{25}-1)}{0.08}"

"P=209,214.02"